OpenStudy (anonymous):

Simplify the expression x radical 64x^3 + 2x radical x + 8 radical 2x

OpenStudy (anonymous):

@asnaseer could ya help

OpenStudy (asnaseer):

ok - first try to simplify each term of this expression

OpenStudy (asnaseer):

lets start with the first term: \(x\sqrt{64x^3}\) do you know what the square root of 64 is?

OpenStudy (anonymous):

8!

OpenStudy (asnaseer):

correct. now when you have an even power inside the square root then you know from one of your previous questions that its square root is just half of the power - remember?

OpenStudy (asnaseer):

e.g. \(\sqrt{m^{16}}=m^8\)

OpenStudy (anonymous):

Yes. But this is odd . 3

OpenStudy (asnaseer):

yup - so we can make use of a trick here...

OpenStudy (asnaseer):

we know we can write \(x^3=x^2\times x\) agreed?

OpenStudy (anonymous):

Yes!!

OpenStudy (asnaseer):

therefore: \(\sqrt{x^3}=\sqrt{x^2\times x}=\sqrt{x^2}\times\sqrt{x}=?\) can you work this out?

OpenStudy (anonymous):

How do you multiply radical x by radical x ^2 ....?

OpenStudy (asnaseer):

lets use some numbers to illustrate this...

OpenStudy (asnaseer):

\[\sqrt{8}=\sqrt{2^3}=\sqrt{2^2\times2}=\sqrt{2^2}\times\sqrt{2}=2\sqrt{2}\]because: \(\sqrt{2^2}=2\)

OpenStudy (asnaseer):

make sense?

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

good, so can you work this one out? \[\sqrt{x^3}=\sqrt{x^2\times x}=\sqrt{x^2}\times\sqrt{x}=?\]

OpenStudy (anonymous):

So just x radical x ?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so, going back to the first term, we get: \(x\sqrt{64x^3}=x\times8\sqrt{x^3}=x\times8x\sqrt{x}=8x^2\sqrt{x}\) agreed?

OpenStudy (anonymous):

Yessir!

OpenStudy (asnaseer):

so now your original question was to simplify:\[x\sqrt{64x^3} + 2x\sqrt{x} + 8\sqrt{2x}\]is this correct?

OpenStudy (anonymous):

Tis correct.

OpenStudy (asnaseer):

so using our last result we get:\[x\sqrt{64x^3} + 2x\sqrt{x} + 8\sqrt{2x}=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2x}\]

OpenStudy (asnaseer):

now notice that \(\sqrt{2x}=\sqrt{2}\times{\sqrt{x}}\), so we can write this as:\[x\sqrt{64x^3} + 2x\sqrt{x} + 8\sqrt{2x}=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2x}\]\[\qquad=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2}\sqrt{x}\]make sense so far?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

good. now can you see any common factors between ALL the terms?

OpenStudy (anonymous):

TWO,! :D

OpenStudy (asnaseer):

yes - that is ONE of the factors. can you spot another one?

OpenStudy (anonymous):

4 or 8...? Idk

OpenStudy (asnaseer):

can you see that every term has \(\sqrt{x}\) in it?

OpenStudy (anonymous):

Yes oh so like 2x?

OpenStudy (asnaseer):

almost - the common factor in every term is \(2\sqrt{x}\)

OpenStudy (asnaseer):

so we can take this common factor out to get:\[x\sqrt{64x^3} + 2x\sqrt{x} + 8\sqrt{2x}=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2x}\]\[\qquad=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2}\sqrt{x}\]\[\qquad=2\sqrt{x}(4x^2+x+4\sqrt{2})\]

OpenStudy (anonymous):

So that is the simplified expression?

OpenStudy (asnaseer):

yes - I cannot see any further simplification for it

OpenStudy (anonymous):

In standard form?

OpenStudy (asnaseer):

what have you been taught in terms of "standard form"?

OpenStudy (asnaseer):

my /guess/ would be that you have been taught that this is the "standard form":\[\qquad=8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2}\sqrt{x}\]

OpenStudy (asnaseer):

in which case you can stop at this step. otherwise you can go one step further and factorise as we did above.

OpenStudy (anonymous):

Well we want the expression simplified and then written in standarn

OpenStudy (anonymous):

Standard form *

OpenStudy (asnaseer):

you write the answer as:\[8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2}\sqrt{x}\]and then add a sentence saying that this can be further factorised to get:\[2\sqrt{x}(4x^2+x+4\sqrt{2})\] that way you have covered all bases. :)

OpenStudy (asnaseer):

*you could write...

OpenStudy (anonymous):

Well the question asks me to state the leading coefficient and idk if it is

OpenStudy (anonymous):

2 radical x or 8x^2 ... @asnaseer

OpenStudy (asnaseer):

coefficients are usually just numbers, so it would be 8 in this case

OpenStudy (asnaseer):

\[8x^2\sqrt{x} + 2x\sqrt{x} + 8\sqrt{2}\sqrt{x}\]

OpenStudy (asnaseer):

hope that made sense?

OpenStudy (asnaseer):

the leading term here is: \(8x^2\sqrt{x}\) the leading coefficient here is: 8